AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 18 of 33  ELOG logo
ID Date Author Subject
  324   Mon Feb 28 07:31:25 2022 JG, AKCurrent status

08:30

Temperature setpoint = 110 k

Rutherford rates shaky, peak max at ~ 3x10^2

Merger OK

Data rate ~ 1500 kb/sec

Ring avg pressure ~ 2.7E-10 mbar

Temperatures OK

System checks OK - wr fail attached

Leakage current ~ 3.22 uA

  323   Mon Feb 28 06:12:58 2022 JG, AK, USCurrent status

07:20

Temperature setpoint = 110 k

Rutherford rates shaky, peak max at ~ 3x10^2

Merger OK

Data rate ~ 1200 kb/sec

Ring avg pressure ~ 2.5E-10 mbar

Temperatures OK

System checks OK - wr fail attached

Leakage current ~ 3.222 uA

  322   Mon Feb 28 05:04:47 2022 JM, OFTemperature cycle to 125K

06:01

Rutherford rate was decreased compared to previous values. Max rate ~ 50 counts for the peak, average ~20 counts. 

Temperature increased to 125 K

No large increase in chamber pressure observed. 

06:25

Temperature setpoint back to 110 K.

Rutherford rate has increased, however is quite variable. Peak height 3*10^1 to 10^2 range. 

06:54

Setpoint = 110 K

Ruthrford rate has stabilised around 7*10^1 to 10^2 peak height.

Merger OK

Data rate ~ 1300 kb/sec

Leakage current ~ 3.22 uA

Temperatures OK

System checks OK

  321   Mon Feb 28 04:14:26 2022 JM, OFCurrent status

05:20

Temperature setpoint = 110 k

Rates are stable however are slightly lower than previous. Min peak height ~40 counts, max ~ 10^2.

Merger OK

Data rate ~ 1200 kb/sec

Ring avg pressure ~ 2.9E-10 mbar

Temperatures OK

System checks OK - wr fail attached

Leakage current ~ 3.222 uA

  320   Mon Feb 28 03:16:37 2022 JM, OFCurrent status

04:24

Temperature setpoint 110 K

Rates appeared stable throughout the hour ~10^2 peak height. Last 5 minutes rutherford peak height appears reduced, cycling from 10^1 to 10^2. If this continues with sustained low density, temperature will be cycled to 125 K and back down.

Merger OK

Data rate = 1300 kb/sec

Ring avg pressure = 3.1E-10 mbar

Temperatures OK

System checks OK - wr fail attached

Leakage current ~ 3.228 uA

04:33

Rutherford rate appears to have increased and stabilised again around 10^2 peak height.

  319   Mon Feb 28 02:16:00 2022 JM, OFCurrent status

03:21

Temperature setpoint 110 K

Rutherford rate stable. Min peak height ~ 10^2 counts, max ~ 3*10^2 counts.

Temperatures OK

System checks OK - wr attached

Merger OK

Data rate ~ 1500 kb/sec

Leakage current ~ 3.228 uA

Ring vacuum stable, Avg = 3.1E-10 mbar

E1 pump current = 2.5 A

  318   Mon Feb 28 01:16:02 2022 JM, OFTemperature decrease 115K to 110K

02:15

Density remains stable at 155 K

Temperature setpoint decreased by 1K. Aim to get to 110 K to see if higher stable density can be achieved. 

Spike in YR pressures observed as temperature was decreased, still look OK. Will leave to settle for a while. T = 114K.

Rates remain stable at ~ 10^2 peak height level.  

02:33

Pressures stabilised

Setpoint to 113 K. 

02:40

No pressure spike observed after decrease

Setpoint to 112 K

02:46

No pressure spike observed after decrease

Setpoint to 111 K

Pressure spike observed  in YR pressures (smaller than last) - give time to recover. 

03:00

Setpoint to 110 K

No pressure spike observed. 

Rutherford rate peak height remains stable at ~ 10^2 . Perhaps peak height minimum is consistently higher than before, difficult to say. 

Attachment 1 - 115 K

Attachment 2 - spike 1

Attachment 3 - spike 2

All pressure stable now 

  317   Mon Feb 28 00:24:05 2022 JM, OFCurrent status

01:23

Temperature setpoint = 115 k

Rutherford rates still appears stable. Max peak ~10^2 counts, min peak ~50 counts per 20 seconds. Greater fluctuations in YR07 plot however do not appear too high, peaks coincide with sustained higher rutherford counts  ~10^2 counts most of the time.

Ring Avg pressure 3.1 E-10 mbar

Ring current ~ 6E7

Temperatures OK

System wide checks OK - wr fails is attached

Merger OK

Data rate ~ 1400 kb/sec

Leakage current ~ 3.23 uA

  316   Sun Feb 27 23:17:33 2022 JM, OFCurrent status

00:20

Temperature setpoint 115 K

Rutherford rate appears fairly stable. Fluctuations from a max peak of 10^2 to a minimum of ~40 peak hieght. Fluctuates remain lower than at 95 K.

Temperatures OK

System checks OK - white rabbit attached.

Leakage current ~ 3.236 uA

Data rate ~1177 kb/sec

Merger OK

Ring vacuum levels appear to have stablised from previous large fluctuations. Avg ring pressure decreased to 3.8E-10 mbar from 7.5E-10 mbar at start of the shift. (attachment 9)

00:51

Rutherford rate still stable. Fluctuations have reduced slightly. Min peak height now ~60 counts, max peak ~10^2 counts. (attachment 10)

Ring vacuum avg = 3.2 E-10 mbar.

E1 pump current ~ 2.4 A

  315   Sun Feb 27 22:03:51 2022 JM, OFCurrent Status

23:00

Target Temperature setpoint 95 K

Current in the ring ~ 5e7 pps

Rutherford rate appears to fluctuate by 2 ordera of magnitude depending on target density. Rate currently ~10^2 counts per 20s. (attachmet 1) can vary to nothing (attachment 6). When consistently low will increase temperature

Effect of target density can be observed in beam current plots. (attachment 7). When rutherford rate is high, beam current tends to decrease over the shot. When the rate is low the current maintains a level throughout. 

Merger OK

Data rate ~ 800 kb/sec. Current run 78

Leakage current ~3.24 uA

Temperatures OK

System Checks OK - white rabbit fail attached

23:36

Rutherford rate consistently low. Maximum rate was just below 10^2 counts, only lasted for one beam shot then low count rates < 10^1 counts.

Setpoint increase to 125 K. 

Increase in pressure across the board in gasjet after increasing temperature which reduces fairly rapidly after 5 minutes. Current in E1 pump reduced from 2.7 A to 2.3 A at 125k. 

Rate at 125K attachment 8. Appears stable.

23:49

Setpoint decrease to 115 K,

23:56

Temperature now stable at 115K. Rate has increased to ~10^2 counts per 20 secs and appears stable (no large fluctuations). Will keep at this level to see if it can be maintained for significant time period.

  314   Sun Feb 27 20:38:06 2022 Rui-Jiu Chen, YuriCurrent status

Ring current ~5.0e7 particles/s

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 1081kbytes/s

SW4

- current plot uploaded. 3.25 uA.

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

  313   Sun Feb 27 20:19:06 2022 Rui-Jiu Chen, Yuriopen the valves of target.
  312   Sun Feb 27 19:01:02 2022 Rui-Jiu Chen, YuriCurrent status

Ring current ~5.96e7 particles/s

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 1062 kbytes/s

SW4

- current plot uploaded. 3.25 uA.

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

  311   Sun Feb 27 18:46:42 2022 Rui-Jiu Chen, YuriChange target temperature

time              Temperature             E1 current(A)                 E1 frequency(Hz)

19:06           170                                 1.2                                 820

19:11           200                                 1.53                               820

19:37           230                                 1.56                               820

19:37           275                                 1.35                               820

19:48           300                                 1.2                                 820

keep this temperature for 15 minues.

20:07 change the temperature from 300 k to 95 k.

21:11          95                                   2.65                               820

 

 

  310   Sun Feb 27 17:00:45 2022 Rui-Jiu Chen, Yuricurrent status

Ring current ~7.76e7 particles/s

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 1664 kbytes/s

SW4

- current plot uploaded. 3.25 uA.

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

  309   Sun Feb 27 16:30:35 2022 Rui-Jiu Chen, YuriChange target temperature again

1. change T from 95 k to 140 k.

2. keep T at 140k for 15 minutes.

3. change T from 140 k to 95 k.

  308   Sun Feb 27 15:41:33 2022 Rui-Jiu Chen, Yuricurrent status

16:38 change target temperature from 95 k to 140 k.

16:48 change target temperature from 140k to 95k.

 

Ring current ~5.62e7 particles/s

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 1190 kbytes/s

SW4

- current plot uploaded. 3.25 uA.

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

  307   Sun Feb 27 15:02:16 2022 Rui-Jiu Chen, YuriChange target temperature

16:00 change temperature from 95 k to 120k.

16:07 change temperature from 120 k to 95k.

 

Ring current ~4.97e7 particles/s

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 1036 kbytes/s

SW4

- current plot uploaded. 3.25 uA.

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

  306   Sun Feb 27 14:05:30 2022 Rui-Jiu, RaganCurrent status

Ring current ~7.61e7 particles/s

SW3

- Merger OK

- Screenshot uploaded

- Data rate : 1312 kbytes/s

SW4

- current plot uploaded. 3.25 uA.

SW2

- T OK

- system wide checks OK

- Check statistics tab ok

- Spectra ok

SW5

4 plots uploaded

  305   Sun Feb 27 13:34:24 2022 RSS TDCurrent status

14:33: 95K-->140K

14:44: 140K --> 95K

Ring current ~6e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.43 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

ELOG V3.1.4-unknown